Hi
The ratio of henry's eggs to Ivan's to Nellie's was 3;5;10 respectively. When nellie gave 56 eggs to be shared between Henry and ivan she found that the 3 had the same number of eggs.
How many eggs did Nellie have left
Thanks

Answer by ikleyn(52878)

.
The ratio of Henry's eggs to Ivan's to Nellie's was 3;5;10 respectively. When Nellie gave 56 eggs to be shared
between Henry and ivan she found that the 3 had the same number of eggs.
How many eggs did Nellie have left
~~~~~~~~~~~~~~


Solution

They had initially/originally 3x, 5x and 10x  eggs, where x is some (unknown) integer number.


In all, they had originally  3x + 5x + 10x = 18x  eggs.


After exchange, they have equal number of eggs, equal to  18x/3 = 6x for everyone.


Hence, Henry obtained 6x-3x = 3x eggs;  Ivan obtained  6x-5x = x eggs from Nellie.


We know that the sum  3x + x = 4x is equal to 56:


    4x = 56.


Hence, x = 56/4 = 14  eggs.


ANSWER.  Nellie has  6x = 6*14 = 84 eggs left.

Solved, answered, explained and completed.